### v1

# Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction

Abstract
Aspect-Opinion Pair Extraction (AOPE) and Aspect Sentiment Triplet Extraction (ASTE) have gained significant attention in natural language processing. However, most existing methods are a pipelined framework, which extracts aspects/opinions and identifies their relations separately, leading to a drawback of error propagation and high time complexity. Towards this problem, we propose a transition-based pipeline to mitigate token-level bias and capture position-aware aspect-opinion relations. With the use of a fused dataset and contrastive learning optimization, our model learns robust action patterns and can optimize separate subtasks jointly, often with linear-time complexity. The results show that our model achieves the best performance on both the ASTE and AOPE tasks, outperforming the state-of-the-art methods by at least 6.98% in the F1 measure. The code is available at https://github.com/Paparare/trans_aste.
Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction
Xinmeng Hou1  Lingyue Fu2  Chenhao Meng2  Hai Hu2thanks: Corresponding author 1Columbia University, New York, NY, USA 2Shanghai Jiao Tong University, Shanghai, China fh2450@tc.columbia.edu fulingyue@sjtu.edu.cn, chenhaomeng@sjtu.edu.cn, hu.hai@sjtu.edu.cn
1 Introduction
Aspect-Based Sentiment Analysis (ABSA) is a fine-grained sentiment analysis task that identifies specific aspects in text and analyzes the sentiments linked to them. It involves subtasks such as: Aspect Extraction (AE): Identifying mentioned aspects. Opinion Extraction (OE): Extracting opinion terms related to aspects. Aspect-Opinion Pair Extraction (AOPE): Extracting aspect-opinion pairs. Aspect-Sentiment Triplet Extraction (ASTE): Identifying aspect, opinion, and sentiment polarity triplets.
Figure 1: For the sentence ’Gourmet food is delicious. Good service, but not so welcoming,’ this example demonstrates the processing steps in both classic and transitional methods for extracting aspect-opinion pairs and tagging sentiment polarity. AE stands for Aspect Extraction, OE for Opinion Extraction, POS represents positive sentiment, and NEG represents negative sentiment.
Classic ABSA approaches often use a pipeline architecture, performing subtasks like aspect extraction (AE) and opinion extraction (OE) separately. For instance, in the sentence: “Gourmet food is delicious. Good service, but not so welcoming.,” AE identifies gourmet food and service as aspects, while OE extracts delicious, good, and not so welcoming as opinions. These outputs are combined to form aspect-opinion pairs, with sentiment tagging assigning polarities to create triplets. Past studies have generally achieved satisfactory performance on aspect and opinion extraction Jiang et al. (2023); Wang et al. (2023); Chakraborty (2024). Adding the according sentiment tags with pair, ASTE is the most integrated task for aspect-based sentiment analysis that has seen the development of diverse models leveraging various methodologies. Pipeline-based approaches, such as Peng-Two-stage Peng et al. (2020), decompose the task into multiple stages for modular refinement. Sequence-to-sequence frameworks like BARTABSA Yan et al. (2021) employ pretrained transformers to generate triplets flexibly. Sequence-tagging methods, including GTS Wu et al. (2020) and JET-BERT Xu et al. (2020), annotate tokens for precise identification of relationships. Machine Reading Comprehension (MRC)-based models, such as COM-MRC Zhai et al. (2022) and Triple-MRC Zou et al. (2024), reframe the task as query answering for efficient extraction. Graph-based approaches such as EMC-GCN Chen et al. (2022), BDTF Chen et al. (2022), and DGCNAP Li et al. (2023) use graph structures to capture semantic and syntactic interactions. Tagging schema-based models, exemplified by STAGE-3D Liang et al. (2023), use hierarchical schemas for multi-level extraction, while lightweight models like MiniConGTS Sun et al. (2024) focus on efficiency with reduced computational costs.
Although previous studies have shown satisfactory results for individual entity extraction, performance on pair and triplet extraction remains suboptimal, with state-of-the-art (SOTA) scores in the 60s and 70s Sun et al. (2024). Two key challenges contribute to this gap: (1) Disconnected Aspect-Opinion Extraction: Opinions are often extracted independently from their corresponding aspects in previous studies Liang et al. (2023); Sun et al. (2024). While relationships can be added as an auxiliary factor to assist pair extraction Liu et al. (2022); Wang et al. (2023), this approach loses critical contextual information by treating aspects and opinions as separate entities. This limits the effectiveness of many token-based extraction methods. (2) Time Complexity with Longer Sequences: Methods using 2D matrix tagging Liang et al. (2023); Sun et al. (2024) to capture relationships between tokens face significant increases in time complexity as the length of the token sequence increases. This computational burden restricts their scalability, especially for longer texts in practical applications.
As an overview, the critical contributions of our study are as follows:
Efficient Transition Pipeline for AOPE:
We propose a transition pipeline for simultaneous aspect-opinion pair extraction, achieving linear time complexity $O(n)$.
Augmented Learning for Model Optimization:
Building on this transitional backbone, we introduce a contrastive-augmented optimization method to enhance model efficiency.
New SOTA Performance on AOPE and ASTE:
Our model achieves new state-of-the-art (SOTA) performance on both AOPE and ASTE tasks.
2 Related Work
2.1 Transition-based Methods
Transition-based approaches are widely used in dependency parsing, leveraging shift-reduce and bidirectional arc actions (left-arc, right-arc) for efficient $O(n)$ parsing Aho and Ullman (1973); Nivre (2003); Cer et al. (2010). These parsers maintain stack, buffer, and arc relations to track transitions and have been extended to tasks like token segmentation Zhang et al. (2016) and argument mining Bao et al. (2021). Beyond parsing, transition-based methods have been applied to structured sentiment analysis, generating graph structures for opinion extraction Fernández-González (2023) and incremental constituency parsing using graph neural networks Yang and Deng (2020). A unifying theory bridges transition-based parsing and sequence labeling for simplified learning Gómez-Rodríguez et al. (2020), while in AMR parsing, structure-aware fine-tuning achieves state-of-the-art results Zhou et al. (2021). Their flexibility and efficiency make them effective across diverse NLP tasks.
2.2 Contrastive-based Optimization
Contrastive learning has demonstrated its effectiveness in various domains, achieving state-of-the-art results in token-independent extraction tasks with models such as MINIconGTS Sun et al. (2024). Recent works have further explored its potential: contrastive learning for prompt-based few-shot language learners enhances generalization through augmented "views" Jian et al. (2022); contrastive learning as goal-conditioned reinforcement learning provides a novel loss function with theoretical guarantees for improved success in goal-directed tasks Eysenbach et al. (2023), and studies analyzing the role of margins reveal critical factors like positive sample emphasis for better generalization Rho et al. (2023). These advances showcase contrastive learning’s flexibility and robustness, which inspired us to integrate it into our transition-based AOPE and ASTE to refine representation learning and improve model performance.
3 Methodology
Action Symbolic Expression
Shift ($SF$) $(\sigma_{0},\,\beta_{0}\mid\beta_{1},E,C,R)\xrightarrow{SH}(\sigma_{0}\mid\sigma_{1},\beta_{1},E,C,R)$
Stop ($ST$) $(\sigma_{0},\O,A,O,R)\xrightarrow{ST}(\O,\O,A,O,R)$
[…]
Left Constituent Removal ($L_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{L_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
Right Constituent Removal ($R_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{R_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
Left-Relation Formation ($LR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},A,O,R)\xrightarrow{LR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{1},O\cup\sigma_{0},R\cup\sigma_{0}\xleftarrow{}\sigma_{1})$
Right-Relation Formation ($RR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},E,C,R)\xrightarrow{RR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{0},O\cup\sigma_{0},R\cup\sigma_{0}\xrightarrow{}\sigma_{1})$
Table 1: Symbolic Expressions for the Proposed Actions.
In this section, we introduce a novel paradigm that integrates the extraction of aspect-opinion pairs into a process reminiscent of constructing a parsing-directed graph. This model progressively constructs and annotates aspect-oriented relations using input sequences enriched with contextual features. Unlike the classic approach, which employs three sets to record action histories, we utilize five sets, including stack, buffer, aspect, and pair, to show how aspect and opinion are extracted sitimulately.
3.1 Transitional Operations and State Change
Phrase Action Stack ($\sigma$) Buffer ($\beta$) Aspect Opinion Pair
[] [$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
1 $SF$ [$\sigma_{1}$] [$\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
2 $SF$ [$\sigma_{1}$, $\sigma_{2}$] [$\beta_{3}$, $\beta_{4}$]
3 $M$ [$\sigma_{1\&2}$] [$\beta_{3}$, $\beta_{4}$]
4 $SF$ [$\sigma_{1\&2}$, $\sigma_{3}$] [$\beta_{4}$]
5 $R_{n}$ [$\sigma_{1\&2}$] [$\beta_{4}$]
6 $SF$ [$\sigma_{1\&2}$, $\sigma_{4}$] []
7 $RR$ [$\sigma_{1\&2}$, $\sigma_{4}$] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
Table 2: State changes for "Gourmet food is delicious" using symbolic representation. Here, $\sigma_{1}$ corresponds to "Gourmet", $\sigma_{2}$ to "food", $\sigma_{3}$ to "is", and $\sigma_{4}$ to "delicious". Similarly, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ correspond to tokens in the buffer in sequence.
The construction of relationships between phrases is achieved through systematic transition actions, conceptualized as directed edges linking two nodes (tokens), $N_{1}$ and $N_{2}$. A relationship transitioning $N_{1}$ to $N_{2}$ under relation $l_{1}$ is denoted as $RR=N_{1}\xrightarrow{l_{1}}N_{2}$ or $LR=N_{1}\xleftarrow{l_{1}}N_{2}$, representing aspect-oriented direction. To interpret causal links, $l_{1}$ is defined as $l_{L}$ or $l_{R}$ for bidirectional relations. Opinion constituents ($E_{1}$) may precede or follow aspect constituents, which can span multiple tokens (e.g., “gourmet food,” “not bad”). This requires additional steps to consolidate multi-token constituents. For the ASTE task, we define seven distinct transition actions that combine token retrieval, termination, and token merging.
The action framework aims to prevent information leakage and ensure flexibility in relation arching. Actions are recorded as tuples $T=(\sigma,\beta,A,O,R)$, representing the stack, buffer, aspect, opinion, and relation. The ASTE task includes default actions, which are universal and required for all data, primary actions that handle merging or removal, and secondary actions for relation arching, which depend on the outcomes of the primary actions. Verbal and symbolic representations of state transitions for each action are provided below.
Default Actions:
  1. 1.
    Shift ($SF$) moves a token from the tokenized stack into the buffer for further processing.
  2. 2.
    Stop ($ST$) halts the process when only one token remains in the buffer, and the stack is empty.
Primary Actions:
  1. 1.
    Merge ($M$) combines multiple tokens in the buffer into a single compound target.
  2. 2.
    Left Constituent Removal ($L_{n}$) removes the left constituent from the buffer.
  3. 3.
    Right Constituent Removal ($R_{n}$) removes the right constituent from the buffer.
Secondary Actions:
  1. 1.
    Left-Relation Formation ($LR$) creates a relation from the right aspect constituent to the left opinion constituent.
  2. 2.
    Right-Relation Formation ($RR$) creates a relation from the left aspect constituent to the right opinion constituent.
Table 1 provides a symbolic illustration of how the symbolic state is constructed and utilized. Take the sentence "Gourmet food is delicious" as an example. Table 2 demonstrates the process of moving tokens from the buffer to the stack, deciding whether they should be merged into a single entity or removed, and finally evaluating them for relation formation. It is important to note that the set of actions shown in the figure is not the only way to extract the "Gourmet food" and "delicious" aspect-opinion pair. The alternative solution for parsing the sentence is shown in Appendix 5.
3.2 Trans-AOPE State Representation
Figure 2: The complete process of the transition-based model is illustrated. Purple highlights represent the transition-based pair extraction, while orange indicates the final step of sentiment tagging.
The model we propose consists of two core stages: pair extraction with a designed transitional action slot (in purple) and pair-based sentiment tagging (in orange), as illustrated in Figure 2.
In the first stage, the input, denoted as $I^{n}_{1}=(t_{1},t_{2},\ldots,t_{n})$, is a sequence of tokens. The output is a sequence of actions, represented as $A^{m}_{1}=(a_{1},a_{2},\ldots,a_{m})$. This process can be conceptualized as a search for the optimal action sequence, $A^{*}$, given the input sequence $I^{n}_{1}$. At each step $n$, the model predicts the next action based on the current system state, $S$, and the sequence of prior actions, $A^{n-1}_{1}$ The updated system state, $S_{n+1}$ , is determined by the specific action $a_{t}$. We define $r_{n}$ as a symbolic representation for calculating the probability of the action $a_{n}$ at step $n$. This probability is computed as follows:
$p(a_{n}|r_{n})=\frac{\exp(w_{a_{n}}^{\top}r_{n}+b_{a_{n}})}{\sum\nolimits_{a^{\prime}\in\mathcal{A}(S)}{\exp(w_{a^{\prime}}^{\top}r_{n}+b_{a^{\prime}})}}$ (1)
Here, $w_{a}$ is a learnable parameter vector, and $b_{a}$ is a bias term. The set $\mathcal{A}(S)$ represents the legal actions available given the current parser state. The overall optimization objective for the model is defined as:
$(A^{*},S^{*})$ $=\underset{{}_{A,S}}{\mathrm{argmax}}\prod_{n}p(a_{n},S_{n+1}|A^{n-1}_{1},S_{n})$ (2)
$=\underset{{}_{A,S}}{\mathrm{argmax}}\prod_{n}p(a_{n}|r_{n})$
In this stage, the ASTE task is incorporated into a transition-based action prediction task. In order to decode efficiently, the action boasting the maximum probability is chosen through a greedy approach until the parsing procedure reaches a point of termination. This model thus forms the basis of an efficient transition-based prediction system that utilizes a representation of the current system state and action history to generate the most likely subsequent action at each step in a sequence. By integrating these concepts, the model successfully carries out parsing tasks while avoiding information leakage and ensuring flexibility in the relation arching process.
3.3 Transition Implementation with Neural Model
This section outlines the transition-based parsing process using neural models. RoBERTa Liu et al. (2019) encodes textual phrases, while UniLSTM Hochreiter and Schmidhuber (1997) and BiLSTM Graves and Schmidhuber (2005) model dynamic transitions. The parser state is initialized and evolves through a sequence of actions with distributed embeddings and LSTMs representing the state at each step. At the end stage, a Multi-Layer Perceptron (MLP) is employed to classify the sentiment associated with each pair or triplet based on the evolved parser state.
Token representations
Consider the process of parsing a text $d^{n}_{1}$ = ($p_{1}$, $p_{2}$, $\ldots$, $p_{n}$) , consisting of $n$ phrases. Each phrase $p_{i}$ = ($w_{i1}$, $w_{i2}$, $\ldots$, $w_{il}$) contains $l$ tokens. A phrase can be represented as a sequence $x_{i}$ = ($\mathrm{[CLS]}$, $t_{i1}$, $\ldots$, $t_{il}$, $\mathrm{[SEP]}$), where $\mathrm{[CLS]}$ is a special classification token whose final hidden state serves as the aggregate sequence feature, and $\mathrm{[SEP]}$ is a separator token. The hidden representation of each phrase is computed as $h_{p_{i}}$ = $\mathrm{RoBERTa}$($x_{i}$) $\in\mathbb{R}^{d_{b}\times|l_{i}|}$, where $d_{b}$ is the hidden dimension size, and $|l_{i}|$ is the length of the sequence $x_{i}$. Finally, the entire text $d^{n}_{1}$ is represented as a list of tokens: $h_{d}$ = [$h_{p_{1}}$, $h_{p_{2}}$, $\ldots$, $h_{p_{n}}$].
State Initialization
At the start of the parsing process, the parser’s state is initialized as $(\beta=\varnothing,\sigma=[1,2,\ldots,n],E=\varnothing,C=\varnothing,R=\varnothing)$, where $\sigma$ is the stack, $\beta$ is the buffer, and $E$, $C$, and $R$ are empty sets representing different outputs. The state evolves through a sequence of actions, progressively consuming elements from the buffer $\beta$ and constructing the output. This process continues until the parser reaches its terminal state when there is only one token left in buffer, represented as $(\beta=[SEP],\sigma=\varnothing,E,C,R)$.
Step-by-Step Parser State Representation.
For the action sequence, each action $a$ is mapped to a distributed representation $e_{a}$ through a lookup table $E_{a}$. An unidirectional LSTM is then utilized to capture the complete history of actions in a left-to-right manner at each step $t$:
$\alpha_{t}=\text{LSTM}_{a}(a_{0},a_{1},\ldots,a_{t-1},a_{t})$ (3)
Upon generation of a new action $a_{t}$, its corresponding embedding $e_{a_{t}}$ is integrated into the rightmost position of LSTMa. To further refine the representation of the pair $(\sigma_{1},\sigma_{0})$, their relative positional distance $d$ is also encoded as an embedding $e_{d}$ from a lookup table $E_{d}$. The composite representation of the parser state at step $t$ encompasses these varied features.
The parser state is represented as a triple $(\beta_{s},\sigma_{s},A_{t})$, where $\sigma_{s}$ denotes the stack sequence $(\sigma_{0},\sigma_{1},\ldots,\sigma_{n})$, $\beta_{s}$ represents the buffer sequence $(\beta_{0},\beta_{1},\ldots,\beta_{n})$, and $A_{t}$ encapsulates the action history $(a_{0},a_{1},\ldots,a_{t-1},a_{t})$. The stack ($\sigma_{n}$) and buffer ($\beta_{n}$) are encoded using bidirectional LSTMs as follows:
$=\text{BiLSTM}((\sigma_{n},\beta_{0}),(\sigma_{n-1},\beta_{1}),$ (4)
$\ldots,(\sigma_{0},\beta_{n}))$
Here, $s_{t}$ and $b_{t}$ are the output feature representations of the stack and buffer, respectively. Each of these representations consists of forward and backward components: $\sigma_{t}=(\overrightarrow{\sigma_{t}},\overleftarrow{\sigma_{t}})$ and $\beta_{t}=(\overrightarrow{\beta_{t}},\overleftarrow{\beta_{t}})$. The forward and backward components are matrices in $\mathbb{R}^{d_{l}\times|\sigma_{t}|}$ and $\mathbb{R}^{d_{l}\times|\beta_{t}|}$, respectively, where $d_{l}$ is the hidden dimension size of the LSTM, and $|\sigma_{t}|$, $|\beta_{t}|$ are the lengths of the sequences $\sigma_{t}$ and $\beta_{t}$.
3.4 Optimization Implementation
We compare two optimization strategies: regular optimization using Cross-Entropy Loss and contrastive-based optimization, which aligns predicted and true action embeddings. A weight study will evaluate the impact of the positioning of two components in augmented optimization on model performance.
3.4.1 Base Optimization
Both action and sentiment classification tasks are optimized using the Cross-Entropy Loss as base optimization, defined as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}y_{i}^{j}\log(p_{i}^{j}),$ (5)
where $N$ is the number of samples, $M$ is the number of classes (either action or sentiment classes), $y_{i}^{j}$ is a binary indicator (0 or 1) indicating whether class $j$ is the correct class for sample $i$, and $p_{i}^{j}$ is the predicted probability for class $j$ for sample $i$. The total loss is the sum of the losses for both tasks:
$\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{sentiment}}.$ (6)
3.4.2 Contrastive-based Optimization
Given the predicted action logits $\mathbf{A}_{\text{logits}}$ and the true action labels $\mathbf{A}_{\text{true}}$, the predicted actions are obtained by applying a softmax function followed by an argmax:
$\mathbf{A}_{\text{pred}}=\text{argmax}(\text{softmax}(\mathbf{A}_{\text{logits}}))$
The predicted and true actions are then embedded as follows:
$\mathbf{E}_{\text{pred}}=\text{Embed}(\mathbf{A}_{\text{pred}}),\quad\mathbf{E}_{\text{true}}=\text{Embed}(\mathbf{A}_{\text{true}})$
For each sample $i$, compute the cosine similarity between $\mathbf{E}_{\text{pred}}^{(i)}$ and all $\mathbf{E}_{\text{true}}^{(j)}$ to form a similarity matrix. Using this matrix, the diagonal elements represent positive pairs, while the non-diagonal elements represent negative pairs.
Next, calculate the positive and negative cosine similarities, and apply the exponential function:
$e_{\text{pos}}=\exp(\text{cos\_sim}(\mathbf{E}_{\text{pred}},\mathbf{E}_{\text{true}})\cdot\mathbf{M}_{\text{pos}})$ (7)
$e_{\text{all}}=\exp(\text{cos\_sim}(\mathbf{E}_{\text{pred}},\mathbf{E}_{\text{true}}))$ (8)
where, cos_sim denote cosine similarity function $\mathbf{M}_{\text{pos}}$ is a mask that selects only positive pairs (diagonal entries).
Now compute the contrastive loss as:
$\mathcal{L}_{\text{con}}=-\frac{1}{N}\sum_{j=1}^{N}\log\left(\frac{\exp_{\text{pos}}^{(j)}}{\exp_{\text{all}}^{(j)}}\right)$ (9)
Finally, the total loss is computed as:
$\mathcal{L}_{\text{total}}=\omega_{1}\cdot\mathcal{L}_{\text{base}}+\omega_{2}\cdot\mathcal{L}_{\text{con}}$ (10)
where $\omega$ are weighting factors, and $\mathcal{L}_{\text{base}}$ represents any regularization loss added to improve model generalization.
4 Experiments
4.1 Dataset
We used well-established benchmarks in Sentiment Analysis: SemEval 2014 (14lap, 14res) Pontiki et al. (2014), SemEval 2015 (15res) Pontiki et al. (2015), and SemEval 2016 (16res) Pontiki et al. (2016). These datasets, originally curated for the Semantic Evaluation (SemEval) competitions, are widely used for aspect-based sentiment analysis and extraction tasks Xu et al. (2021).
We employed a data processing framework to construct sentiment-enhanced dependency graphs for ASTE datasets. This framework utilizes SpaCy Honnibal et al. (2020) to extract syntactic dependencies and integrates sentiment scores from SenticNet Cambria et al. (2017) to generate weighted adjacency matrices. These enriched graph representations aim to expand the feature space and reduce biased learning. Along with tokenized sentences and their corresponding aspect-opinion-sentiment triplets, the processed data supports downstream tasks, including model training and evaluation.
Figure 3: F1 Score over epochs for the model trained on 14res, 14lap, 15res, and 16res datasets and tested on the 15res dataset. The results indicate that training on datasets other than 15res, such as 14res can yield an even better performance, which suggests the presence of richer action patterns in other datasets for the model to learn.
We evaluated the conformity of our transition operations in extracting aspects and opinions, with action evaluation results listed in Appendix A.3. Since our approach predicts actions rather than tokens, we assessed how well the model learns action patterns across individual datasets. For instance, as shown in Figure 3, the model did not perform best on the 15res test set when trained on the 15res training set. This suggests that other datasets, such as 14res, may contain richer action patterns for the model to learn. Consequently, we believe that training on a fused dataset provides a more diverse context, enhancing the model’s ability to learn these actions, which we then evaluated on separate test sets.
Figure 4: Growth on different weights for AOPE task on 14lap testing set.
Figure 5: Growth on different weights for ASTE task on 14lap testing set.
4.2 Baselines
We use two baselines: published scores from prior studies and previous SOTA models run on the fused dataset. Regular baselines include classic attention-based methods (e.g., CMLA Wang et al. (2017)), pipeline approaches (Peng-Two-stage Peng et al. (2020)), sequence-to-sequence frameworks (BARTABSA Yan et al. (2021)), sequence-tagging methods (GTS Wu et al. (2020), JET-BERT Xu et al. (2020)), MRC-based models (COM-MRC Zhai et al. (2022), Triple-MRC Zou et al. (2024)), graph-based approaches (EMC-GCN Chen et al. (2022), DGCNAP Li et al. (2023)), and the SOTA MiniConGTS Sun et al. (2024).
In addition to baselines on separate datasets, we selected two state-of-the-art models with publicly available and runnable code: the most recent model, MiniConGTS Sun et al. (2024), and the earlier state-of-the-art for pair extraction, BARTABSA Yan et al. (2021). We used their original configurations and data processing pipelines, modifying only the training data to include all four training sets from 14lap, 14res, 15res, and 16res. The models were then separately evaluated on the four corresponding test sets.
Models 14Res 14Lap 15res 16res
P R F1 P R F1 P R F1 P R F1
CMLA+Wang et al. (2017) $\diamond$ - - 48.95 - - 44.10 - - 44.60 - - 50.00
Peng-two-stage Peng et al. (2020) $\diamond$ - - 56.10 - - 53.85 - - 56.23 - - 60.04
Dual-MRC Mao et al. (2021) $\diamond$ - - 74.93 - - 63.37 - - 64.97 - - 75.71
SpanMlt Zhao et al. (2020) $\bullet$ - - 75.60 - - 68.66 - - 64.68 - - 71.78
BARTABSA Yan et al. (2021) $\diamond$ - - 77.68 - - 66.11 - - 67.98 - - 77.38
MiniConGTS Sun et al. (2024) $\star$ - - 79.60 - - 73.23 - - 73.87 - - 76.29
SOTAs on fused dataset
BARTABSA-fused 75.40 76.76 76.07 72.36 63.40 67.59 88.85 92.77 90.77 87.06 87.32 87.19
MiniConGTS-fused 81.06 78.82 79.92 75.62 71.61 73.56 93.63 92.72 93.17 78.80 81.82 80.28
Ours on fused dataset
Trans-AOPE ($w_{1}$:$w_{2}$ = 1:1) 91.88 80.83 86.00 85.87 74.68 79.89 94.88 88.25 91.45 94.63 89.10 91.80
Trans-AOPE ($w_{1}$:$w_{2}$ = 1:0) 93.80 80.39 86.58 88.60 74.26 80.80 95.07 88.95 91.91 93.10 89.01 91.01
Table 3: Comparison of different models on multiple datasets for AOPE task. Recall and precision values are omitted where they are not provided. Former best scores are underlined, and current best scores are bold. $\diamond$ are retrieved from Yan et al., 2021. $\bullet$ is retrieved from Zhao et al., 2020, and $\star$ are retrieved from Sun et al., 2024
Models 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
Peng-Two-stage Peng et al. (2020) $\star$ 43.24 63.66 51.46 38.87 50.38 42.87 48.07 57.51 52.32 46.96 64.24 54.21
BARTABSA Yan et al. (2021) $\diamond$ 65.52 64.99 65.25 61.41 56.19 58.69 59.14 59.38 59.26 66.60 68.68 67.62
GTS Wu et al. (2020) $\star$ 67.76 67.29 67.50 57.82 51.32 54.36 62.59 57.94 60.15 66.08 66.91 67.93
JET-BERT Xu et al. (2020) $\star$ 70.56 55.94 62.40 55.39 43.57 51.04 64.45 51.96 57.53 70.42 58.37 63.83
COM-MRC Zhai et al. (2022) $\star$ 75.46 68.91 72.01 58.15 60.17 61.17 68.35 61.24 64.53 71.55 71.59 71.57
EMC-GCN Chen et al. (2022) $\star$ 71.21 72.39 71.78 61.70 56.26 58.81 61.54 62.47 61.93 65.62 71.30 68.33
BDTF Chen et al. (2022) $\star$ 75.53 73.24 74.35 68.94 55.97 61.74 68.76 63.71 66.12 71.44 73.13 72.27
DGCNAP Li et al. (2023) $\star$ 72.90 68.69 70.72 62.02 53.79 57.57 62.23 60.21 61.19 69.75 69.44 69.58
Triple-MRC Zou et al. (2024) $\star$ - - 72.45 - - 60.72 - - 62.86 - - 68.65
STAGE-3D Liang et al. (2023) $\star$ 78.58 69.58 73.76 71.98 53.86 61.58 73.63 57.90 64.79 76.67 70.12 73.24
MiniConGTS Sun et al. (2024) $\star$ 76.10 75.08 75.59 66.82 60.68 63.61 66.50 63.86 65.15 75.52 74.14 74.83
SOTAs on fused dataset
BARTABSA-fused 71.05 72.33 71.68 63.66 56.01 59.59 88.98 92.77 90.77 83.82 84.07 83.95
MiniConGTS-fused 73.89 72.84 73.36 68.26 57.42 62.37 90.64 89.32 89.98 72.73 73.21 72.97
Ours
Trans-ASTE ($w_{1}$:$w_{2}$ = 1:1) 89.30 77.10 82.76 83.81 69.64 76.07 93.80 86.71 90.11 92.18 86.81 89.42
Trans-ASTE ($w_{1}$:$w_{2}$ = 1:0) 91.70 78.24 84.44 84.63 70.59 76.98 94.31 88.11 91.11 92.78 87.58 90.11
Table 4: Comparison of different models on multiple datasets for ASTE task. Former best scores are underlined, and current best scores are bold. $\star$ are retrieved from Sun et al., 2024, and $\diamond$ is retrieved from Yan et al., 2021.
4.3 Results and Analysis
4.3.1 Study on Augmented Optimization
To assess the impact of weight compositions in Equation 10, we tested $(w_{1},w_{2})\in\{(1:0),(1:1),(1:10),(0:1),(10:1)\}$, where $w_{1}$ and $w_{2}$ are the weights for base and contrastive-based optimization, respectively, using a batch size of 4. Results for AOPE and ASTE are presented in Figures 5 and 5. For AOPE, contrastive-heavy weights ($w_{1}:w_{2}=1:1$ and $w_{1}:w_{2}=1:10$) performed well during early epochs (1–50), demonstrating the early advantages of contrastive optimization. After epoch 50, base ($w_{1}:w_{2}=1:0$) and balanced ($w_{1}:w_{2}=1:1$) optimizations provided greater stability and better final performance. In ASTE, base and balanced optimizations consistently outperformed others, indicating that ASTE primarily benefits from foundational learning, with limited advantages from contrastive optimization.
Similar trends were observed across other datasets, as shown in Appendix A.4, highlighting the broader applicability of these findings. Although contrastive-enhanced optimization does not deliver the best overall performance, it effectively improves efficiency in pair extraction tasks.
4.3.2 Comparison with Baselines
On AOPE
[…]
On ASTE
Results for ASTE are listed in Table 4. For the 14Res dataset, Trans-ASTE achieves the highest F1 score (84.44%), representing an APG of 12.76% over BARTABSA-fused. Similarly, it outperforms MiniConGTS-fused with an APG of 11.08%. In the 14Lap dataset, Trans-ASTE) achieves an APG value of 17.39% in F1 compared to BARTABSA-fused, and 14.61% over MiniConGTS-fused. For the 15Res dataset, Trans-ASTE) maintains a high level of performance, with an F1 score (91.11%) slightly leading both baselines. The overall F1 improvement is 0.34% compared to BARTABSA-fused and 1.13% compared to MiniConGTS-fused. The 16Res dataset also sees consistently high performance from Trans-ASTE), with the highest F1 score (90.11%), representing an APG of 6.16% over BARTABSA-fused.
Overall, combining datasets for all-for-one training does not always result in performance improvements across all datasets for SOTA models. However, when data is fused for our model, it consistently achieves performance gains across datasets, with significant improvements observed in 14Res, 14Lap, and 16Res. That said, the slightly lower performance on the 15Res AOPE task highlights potential areas for further optimization in specific scenarios.
5 Conclusion
In this paper, we present an efficient transition pipeline for aspect-opinion pair extraction with linear time complexity $O(n)$, enhanced by a contrastive learning-based optimization method. This approach eliminates the need to directly identify and extract individual tokens, avoiding token-level bias. It can be trained on datasets with the most comprehensive coverage of the designed actions, resulting in significant performance improvements across various datasets. Specifically, training our model on a well-covered fused dataset enables it to learn robust action combination patterns, leading to superior performance on all datasets. Our model surpasses SOTA models on the same fused dataset, setting new benchmarks for AOPE and ASTE tasks. This approach not only enhances model efficiency but also demonstrates its effectiveness and versatility in improving performance.
Limitation
The primary limitation of the transition-based pipeline lies in its reliance on relatively large and diverse datasets to effectively learn action patterns. As evidenced by the findings in Figure 3, the model achieved better performance on the 15res test set when trained on datasets like 14res and 16res, which appear to provide richer and more diverse action patterns compared to the 15res training set. This observation highlights the inadequacy of relying solely on a single dataset, as it may lack the breadth and complexity needed to capture the full spectrum of action patterns required for robust model generalization. Without access to such diverse training data, the model’s ability to adapt to nuanced patterns or unseen contexts in the test data may be significantly compromised, emphasizing the importance of training on a fused or broader dataset to enhance its overall effectiveness.
Appendix A Appendix
A.1 Alternative Solution with State Changes
Since stack can hold more than two tokens, "is" can be moved to stack ($\beta_{3}\rightarrow\sigma_{3}$) before "Gourmet" and "food" merge ($[\sigma_{1},\sigma_{2},\sigma_{3}]\rightarrow[\sigma_{1\&2},\sigma_{3}]$). The detailed process is present in Table 5.
Phrase Action Stack ($\sigma$) Buffer ($\beta$) Aspect Opinion Pair
[] [$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
1 $SF$ [$\sigma_{1}$] [$\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
2 $SF$ [$\sigma_{1}$, $\sigma_{2}$] [$\beta_{3}$, $\beta_{4}$]
3 $SF$ [$\sigma_{1}$, $\sigma_{2}$, $\beta_{3}$] [$\beta_{4}$]
4 $M$ [$\sigma_{1\&2}$, $\beta_{3}$] [$\beta_{4}$]
5 $R_{n}$ [$\sigma_{1\&2}$] [$\beta_{4}$]
6 $SF$ [$\sigma_{1\&2}$, $\sigma_{4}$] []
7 $RR$ [$\sigma_{1\&2}$, $\sigma_{4}$] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
Table 5: State changes for "Gourmet food is delicious" using symbolic representation. Here, $\sigma_{1}$ corresponds to "Gourmet", $\sigma_{2}$ to "food", $\sigma_{3}$ to "is", and $\sigma_{4}$ to "delicious". Similarly, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ correspond to tokens in the buffer in sequence.
A.2 Algorithm for Transition Actions
In Algorithm 1, while the core of the algorithm focuses on the action operations, the configuration also plays a vital role in its functionality.
The algorithm begins with a buffer, a list of input tokens representing the text to be parsed. These tokens are consumed during the parsing process and form the basis for token pair relationships. A critical component of the algorithm is the pair2sent mapping, which associates token pairs with their corresponding sentiment labels, allowing the algorithm to enrich the output with sentiment information, even for token pairs not explicitly provided in the target data (defaulting to a neutral label). Then, each parsing action (e.g., Shift, Left Relation, Right Relation) is mapped to a unique numeric ID. This mapping enables a compact representation of the sequence of parsing actions, which is useful for downstream tasks like training or evaluation. Throughout the parsing process, the algorithm records intermediate states (stack, buffer, actions, and sentiment labels) to maintain a complete history of its operations. Simultaneously, valid token pairs are stored in pairs, capturing the structured relationships in the text.
Algorithm 1 Argument-Opinion Pair Parsing Algorithm
1:buffer (list of input tokens), pair2sent (mapping of token pairs to sentiment labels), action2id (mapping of parsing actions to numeric IDs)
2:states (list of recorded parsing states), pairs (list of valid token pairs)
3:Initialize empty stack, actions, states, and pairs. $\triangleright$ Main processing loop: parse tokens from the buffer
4:while buffer is not empty do
5:     if len(stack) < 2 then
6:         stack.append(buffer.pop(0)) $\triangleright$ Apply $SH$: Shift
7:         Append $SF$ to actions.
8:         continue
9:     end if $\triangleright$ Form a pair from the top two elements in stack
10:     stackPair $\leftarrow$ (stack[-2], stack[-1])
11:     currentAction $\leftarrow$ getAction(stackPair, targetPairs)
12:     Append currentAction to actions. $\triangleright$ Handle relation formation actions ($LR$ and $RR$)
13:     if currentAction == $LR$ then
14:         Add stackPair to pairs.
15:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
16:         Apply stack[-2] and stack[-1] form a left-relation pair.
17:     else if currentAction == $RR$ then
18:         Add stackPair to pairs.
19:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
20:         Apply stack[-1] and stack[-2] form a right-relation pair.
21:     else
22:         sentimentLabel $\leftarrow 3$triangleright Default sentiment label.
23:     end if $\triangleright$ Record the current parsing state
24:     Record (stack, buffer, actions, sentimentLabel) in states. $\triangleright$ Modify the stack based on currentAction
25:     if currentAction == $L_{n}$ then
26:         Remove stack[-2] $\triangleright$ Apply $L_{n}$: Left Constituent Removal
27:     else if currentAction == $R_{n}$ then
28:         Remove stack[-1] $\triangleright$ Apply $R_{n}$: Right Constituent Removal
29:     else if currentAction == $M$ then
30:         Merge stack[-2] and stack[-1] $\triangleright$ Apply $M$: Merge
31:     end if
32:end while $\triangleright$ Finalize by reducing the remaining elements in the stack
33:while len(stack) > 1 do
34:     Remove the top element from stack $\triangleright$ Apply $ST$: Stop
35:end while $\triangleright$ Convert actions to numeric IDs for output
36:Map each action in states to its ID using action2id. return states, pairs.
A.3 Action Coverage on Different Datasets
14lap 14res 15res 16res fused
train (Precision) 99.99 99.99 99.99 99.99 99.99
train (Recall) 85.51 86.63 92.98 93.04 88.73
train (F1) 92.19 92.84 96.36 96.40 94.03
dev (Precision) 99.99 99.99 99.99 99.99 99.99
dev (Recall) 85.97 86.89 94.92 93.04 89.88
dev (F1) 92.46 92.99 97.39 96.40 94.67
test (Precision) 99.99 99.99 99.99 99.99 99.99
test (Recall) 86.66 87.54 91.47 92.42 89.41
test (F1) 92.85 93.36 95.54 96.06 94.41
total (Precision) 99.99 99.99 99.99 99.99 99.99
total (Recall) 86.02 87.03 92.56 92.79 89.00
total (F1) 92.48 93.06 96.14 96.26 94.18
Table 6: Transposed Transitional Action Coverage for 14Lap, 14Res, 15Res, 16Res, and fused Datasets.
The coverage of actions, as shown in Table 6, reveals some differences in the recall values across data sets. The 14Lap and 14Res datasets exhibit lower recall compared to the 15Res and 16Res datasets. This trend is observed across training, development, and testing phases, with recall values for 14 datasets remaining around 85-87%, while 15Res and 16Res achieve significantly higher recalls, exceeding 91% in most cases. The fused dataset strikes a balance with intermediate recall values, reflecting the integration of characteristics from all datasets. This highlights a relative weakness in capturing relevant actions in 14 datasets compared to 15 and 16 datasets.
A.4 Growth with Different Optimization Weight
Figures 7 and 7, 9 and 9, and 11 and 11 depict the performance growth with varying weights applied to the 14res, 15res, and 16res datasets, respectively. These figures highlight the influence of weight adjustments on model accuracy and efficiency in ASTE tasks. By examining the trends across these datasets, we demonstrate the early benefits of augmented optimization in enhancing performance.
Figure 6: Growth on different weights for AOPE task on 14res testing set.
Figure 7: Growth on different weights for ASTE task on 14res testing set.
Figure 8: Growth on different weights for AOPE task on 15res testing set.
Figure 9: Growth on different weights for ASTE task on 15res testing set.
Figure 10: Growth on different weights for AOPE task on 16res testing set.
Figure 11: Growth on different weights for ASTE task on 16res testing set.
A.5 Growth Comparison for AOPE and ASTE
The growth comparison for AOPE and ASTE is illustrated in Figures 13, 13, 15, and 15, corresponding to the 14res, 14lap, 15res, and 16res datasets, respectively. While ASTE generally follows the trends observed in AOPE, the fluctuations in ASTE are more extreme. These figures emphasize the sensitivity of ASTE to variations in model weights or dataset conditions, which results in sharper peaks and deeper troughs compared to the relatively smoother patterns seen in AOPE. This highlights the unique challenges posed by the ASTE task across different datasets.
Figure 12: Growth Comparison for AOPE and ASTE on 14res Testing set.
Figure 13: Growth Comparison for AOPE and ASTE on 14lap Testing set.
Figure 14: Growth Comparison for AOPE and ASTE on 15res Testing set.
Figure 15: Growth Comparison for AOPE and ASTE on 16res Testing set.

### v2

@@
 Abstract
-Aspect-Opinion Pair Extraction (AOPE) and Aspect Sentiment Triplet Extraction (ASTE) have gained significant attention in natural language processing.
-However, most existing methods are a pipelined framework, which extracts aspects/opinions and identifies their relations separately, leading to a drawback of error propagation and high time complexity.
-Towards this problem, we propose a transition-based pipeline to mitigate token-level bias and capture position-aware aspect-opinion relations.
-With the use of a fused dataset and contrastive learning optimization, our model learns robust action patterns and can optimize separate subtasks jointly, often with linear-time complexity.
-The results show that our model achieves the best performance on both the ASTE and AOPE tasks, outperforming the state-of-the-art methods by at least 6.98% in the F1 measure.
-The code is available at https://github.com/Paparare/trans_aste.
-Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction
-Xinmeng Hou1  Lingyue Fu2  Chenhao Meng2  Hai Hu2thanks: Corresponding author 1Columbia University, New York, NY, USA 2Shanghai Jiao Tong University, Shanghai, China fh2450@tc.columbia.edu fulingyue@sjtu.edu.cn, chenhaomeng@sjtu.edu.cn, hu.hai@sjtu.edu.cn
+Aspect-Opinion Pair Extraction (AOPE) and Aspect Sentiment Triplet Extraction (ASTE) have drawn growing attention in NLP.
+However, most existing approaches extract aspects and opinions independently, optionally adding pairwise relations, often leading to error propagation and high time complexity.
+To address these challenges and being inspired by transition-based dependency parsing, we propose the first transition-based model for AOPE and ASTE that performs aspect and opinion extraction jointly, which also better captures position-aware aspect-opinion relations and mitigates entity-level bias.
+By integrating contrastive-augmented optimization, our model delivers more accurate action predictions and jointly optimizes separate subtasks in linear time.
+Extensive experiments on 4 commonly used ASTE/AOPE datasets show that, while performing worse when trained on a single dataset than some previous models, our model achieves the best performance on both ASTE and AOPE if trained on combined datasets, outperforming the strongest previous models in F1-measures (often by a large margin).
+We hypothesize that this is due to our model’s ability to learn transition actions from multiple datasets and domains.
+Our code is available at https://anonymous.4open.science/r/trans_aste-8FCF.
+Train Once for All: A Transitional Approach for
+Efficient Aspect Sentiment Triplet Extraction
 1 Introduction
-Aspect-Based Sentiment Analysis (ABSA) is a fine-grained sentiment analysis task that identifies specific aspects in text and analyzes the sentiments linked to them.
-It involves subtasks such as: Aspect Extraction (AE): Identifying mentioned aspects.
-Opinion Extraction (OE): Extracting opinion terms related to aspects.
-Aspect-Opinion Pair Extraction (AOPE): Extracting aspect-opinion pairs.
-Aspect-Sentiment Triplet Extraction (ASTE): Identifying aspect, opinion, and sentiment polarity triplets.
-Figure 1: For the sentence ’Gourmet food is delicious.
-Good service, but not so welcoming,’ this example demonstrates the processing steps in both classic and transitional methods for extracting aspect-opinion pairs and tagging sentiment polarity.
-AE stands for Aspect Extraction, OE for Opinion Extraction, POS represents positive sentiment, and NEG represents negative sentiment.
-Classic ABSA approaches often use a pipeline architecture, performing subtasks like aspect extraction (AE) and opinion extraction (OE) separately.
-For instance, in the sentence: “Gourmet food is delicious.
-Good service, but not so welcoming.,” AE identifies gourmet food and service as aspects, while OE extracts delicious, good, and not so welcoming as opinions.
-These outputs are combined to form aspect-opinion pairs, with sentiment tagging assigning polarities to create triplets.
-Past studies have generally achieved satisfactory performance on aspect and opinion extraction Jiang et al.
+Aspect-Based Sentiment Analysis (ABSA) is a fine-grained sentiment analysis task that identifies specific aspects in text and analyzes the sentiments linked to them Hu and Liu (2004); Liu (2012); Wang et al.
+(2024).
+As shown in Figure 1, ABSA involves subtasks such as Aspect Extraction (AE) and Opinion Extraction (OE)—identifying mentioned aspects and their related opinions, or the combination—Aspect-Opinion Pair Extraction.
+Once the aspect and opinion have been extracted, a sentiment is usually computed, and this more complicated task is often referred to as Aspect-Sentiment Triplet Extraction (ASTE).
+Figure 1: Demonstration of the processing steps in both classic and transitional methods for extracting aspect-opinion pairs and tagging sentiment polarity.
+Importantly, our proposed transitional method predicts transition actions, and performs pair extraction after the aspect–opinion relationship has been established, allowing the model to capture contextual relationships more effectively.
+AE stands for Aspect Extraction, OE for Opinion Extraction.
+POS and NEG represent positive and negative sentiments respectively.
+For instance, given the sentence: “Gourmet food is delicious.
+Good service, but not so welcoming”, AE identifies gourmet food and service as aspects, while OE extracts delicious, good, and not so welcoming as opinions.
+These outputs are then combined to form aspect-opinion pairs, with a separate sentiment tagging system assigning polarities to create triplets Jiang et al.
 (2023); Wang et al.
 (2023); Chakraborty (2024).
-Adding the according sentiment tags with pair, ASTE is the most integrated task for aspect-based sentiment analysis that has seen the development of diverse models leveraging various methodologies.
+ASTE is the most integrated task for aspect-based sentiment analysis, for which diverse models leveraging various methodologies have been developed, including pipeline-based approach (Peng et al., 2020), sequence-to-sequence method (Yan et al., 2021), sequence-tagging method (Wu et al., 2020; Xu et al., 2020), to name just a few.
+Despite these efforts and growing interests, the accuracy of recent models remains suboptimal, with the best systems scoring 60% or 70% Sun et al.
+(2024).
+There are two key challenges that hinder performance: (1) Disconnected Aspect-Opinion Extraction: Opinions are often extracted independently from their corresponding aspects Liang et al.
+(2023); Sun et al.
+(2024).
+While positional relationships can be added as an auxiliary factor to assist pair extraction Liu et al.
+(2022); Wang et al.
+(2023), this approach loses critical contextual information by treating aspects and opinions as separate entities.
+This limits the effectiveness of many token-based extraction methods.
+(2) High time complexity with longer sequences: Methods using 2D matrix tagging Liang et al.
+(2023); Sun et al.
+(2024) to capture relationships between tokens face significant increases in time complexity as the length of the token sequence increases.
+This computational burden restricts their scalability, especially for longer texts in practical applications.
+To address these two challenges, we present the first transition-based AOPE system named Trans-AOPE that (1) extracts the Aspect and the Opinion at the same time, and (2) has a time complexity of $O(n)$.
+We also introduce a contrastive-augmented optimization method to enhance model efficiency.
+We conduct experiments on 4 commonly used ABAS datasets, and compare our system with previous models.
+Our results show that Trans-model achieves state-of-the-art performance on all datasets we tested.
+We conduct comprehensive ablation studies to evaluate the contribution of optimization components and perform extensive training on various datasets to identify precisely where our model and baselines derive their learning.
+Our contributions are:
+  • We propose the first transition-based model that extracts aspect-opinion pairs based on relational aspects, rather than using relational factors as supplementary references or confirmation, with linear time complexity.
+  • We experiment with a contrastive-augmented optimization method and find that balanced weighting yields faster, more stable improvements, emerging as the optimal training configuration.
+  • We explore various training strategies and show that our proposed method achieves optimal performance on four datasets when trained on combined training sets, with better cross-dataset generalization.
+Method Approach Time Complexity
+Peng-Two-stage (Peng et al., 2020) Two-Stage Pipeline: entity identification and relation formation $O(n+k^{2})$
+BARTABSA (Yan et al., 2021) Generative-based Aspect-based Sentiment Analysis $O(m\cdot v)$
+GTS (Wu et al., 2020) Grid Matrix-based Tagging $O(n^{2})$
+JET-BERT Xu et al. (2020) Position-Aware Sequence Tagging $O(n)$
+COM-MRC Zhai et al. (2022) Compositional Machine Reading Comprehension $O(r\cdot n^{2}\cdot h)$
+Triple-MRC Zou et al. (2024) Multi-turn Machine Reading Comprehension $O(r\cdot n^{2}\cdot h)$
+EMC-GCN Chen et al. (2022) Multi-channel Graph Convolutional Network $O(m\cdot n^{2}\cdot h)$
+DGCNAP Li et al. (2023) Graph Convolutional Network w/ Affective Knowledge $O(m\cdot n^{2}\cdot h)$
+MiniConGTS Sun et al. (2024) Lightweight Grid Matrix-based Tagging System $O(n^{2})$
+Trans-model (Ours) Transition-based Action Prediction for Simulating Relation Formation and Pair Extraction $O(n)$
+Table 1: An overview of previous methods and models (which will serve as baselines in this study), their approaches, and corresponding time complexities.
+Here, the hidden size for LSTM $d$ is simplified; $n$ is the sequence length; $m$ is the number of graph channels; $v$ is the vocabulary size; $k$ is the number of extracted terms; $r$ is the number of query rounds, and $h$ is the hidden size of the encoder.
+2 Related Work
+Previous methods on ASTE
 Pipeline-based approaches, such as Peng-Two-stage Peng et al.
@@
 (2024) focus on efficiency with reduced computational costs.
-Although previous studies have shown satisfactory results for individual entity extraction, performance on pair and triplet extraction remains suboptimal, with state-of-the-art (SOTA) scores in the 60s and 70s Sun et al.
-(2024).
-Two key challenges contribute to this gap: (1) Disconnected Aspect-Opinion Extraction: Opinions are often extracted independently from their corresponding aspects in previous studies Liang et al.
-(2023); Sun et al.
-(2024).
-While relationships can be added as an auxiliary factor to assist pair extraction Liu et al.
-(2022); Wang et al.
-(2023), this approach loses critical contextual information by treating aspects and opinions as separate entities.
-This limits the effectiveness of many token-based extraction methods.
-(2) Time Complexity with Longer Sequences: Methods using 2D matrix tagging Liang et al.
-(2023); Sun et al.
-(2024) to capture relationships between tokens face significant increases in time complexity as the length of the token sequence increases.
-This computational burden restricts their scalability, especially for longer texts in practical applications.
-As an overview, the critical contributions of our study are as follows:
-Efficient Transition Pipeline for AOPE:
-We propose a transition pipeline for simultaneous aspect-opinion pair extraction, achieving linear time complexity $O(n)$.
-Augmented Learning for Model Optimization:
-Building on this transitional backbone, we introduce a contrastive-augmented optimization method to enhance model efficiency.
-New SOTA Performance on AOPE and ASTE:
-Our model achieves new state-of-the-art (SOTA) performance on both AOPE and ASTE tasks.
-2 Related Work
-2.1 Transition-based Methods
+Table 1 summarizes these baseline methods, along with our proposed model, in terms of their core approaches and time complexities.
+Transition-based Methods in NLP
 Transition-based approaches are widely used in dependency parsing, leveraging shift-reduce and bidirectional arc actions (left-arc, right-arc) for efficient $O(n)$ parsing Aho and Ullman (1973); Nivre (2003); Cer et al.
 (2010).
-These parsers maintain stack, buffer, and arc relations to track transitions and have been extended to tasks like token segmentation Zhang et al.
-(2016) and argument mining Bao et al.
+These parsers maintain stack, buffer, and arc relations to track transitions and then build up dependency relations between tokens.
+Transition-based methods have also been applied to various NLP tasks, including token segmentation Zhang et al.
+(2016), argument mining Bao et al.
+(2021), constituency parsing Yang and Deng (2020), AMR parsing Zhou et al.
+(2021), and sequence labeling Gómez-Rodríguez et al.
+(2020), among others.
+Transition-based methods have been explored in emotion analysis Fan et al.
+(2020); Jian et al.
 (2024).
-Beyond parsing, transition-based methods have been applied to structured sentiment analysis, generating graph structures for opinion extraction Fernández-González (2023) and incremental constituency parsing using graph neural networks Yang and Deng (2020).
-A unifying theory bridges transition-based parsing and sequence labeling for simplified learning Gómez-Rodríguez et al.
-(2020), while in AMR parsing, structure-aware fine-tuning achieves state-of-the-art results Zhou et al.
-(2021).
-Their flexibility and efficiency make them effective across diverse NLP tasks.
-2.2 Contrastive-based Optimization
+In sentiment analysis, however, transition-based models have not been widely adopted.
+One exception is their use in generating graph structures for opinion extraction Fernández-González (2023), although this design relies on graph embeddings and thus results in a time complexity of $O(N^{2})$, with performance that lags behind more recent AOPE and ASTE approaches.
+Contrastive-based Optimization
 Contrastive learning has demonstrated its effectiveness in various domains, achieving state-of-the-art results in token-independent extraction tasks with models such as MiniconGTS Sun et al.
 (2024).
-Recent works have further explored its potential: contrastive learning for prompt-based few-shot language learners enhances generalization through augmented "views" Jian et al.
+Recent works have further explored its potential: contrastive learning for prompt-based few-shot language learners enhances generalization through augmented “view” Jian et al.
 (2022); contrastive learning as goal-conditioned reinforcement learning provides a novel loss function with theoretical guarantees for improved success in goal-directed tasks Eysenbach et al.
@@
 These advances showcase contrastive learning’s flexibility and robustness, which inspired us to integrate it into our transition-based AOPE and ASTE to refine representation learning and improve model performance.
-3 Methodology
 Action Symbolic Expression
-Shift ($SF$) $(\sigma_{0},\,\beta_{0}\mid\beta_{1},E,C,R)\xrightarrow{SH}(\sigma_{0}\mid\sigma_{1},\beta_{1},E,C,R)$
-Stop ($ST$) $(\sigma_{0},\O,A,O,R)\xrightarrow{ST}(\O,\O,A,O,R)$
+Shift ($SF$) $(\sigma_{0},\,\beta_{0}\mid\beta_{1},A,O,R)\xrightarrow{SH}(\sigma_{0}\mid\sigma_{1},\beta_{1},A,O,R)$
+Stop ($ST$) $(\sigma_{0},,A,O,R)\xrightarrow{ST}(,,A,O,R)$
 […]
 Left Constituent Removal ($L_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{L_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
-Right Constituent Removal ($R_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{R_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
+Right Constituent Removal ($R_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{R_{n}}(\sigma_{0},\,\beta_{0},A,O,R)$
 Left-Relation Formation ($LR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},A,O,R)\xrightarrow{LR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{1},O\cup\sigma_{0},R\cup\sigma_{0}\xleftarrow{}\sigma_{1})$
-Right-Relation Formation ($RR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},E,C,R)\xrightarrow{RR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{0},O\cup\sigma_{0},R\cup\sigma_{0}\xrightarrow{}\sigma_{1})$
+Right-Relation Formation ($RR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},A,O,R)\xrightarrow{RR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{0},O\cup\sigma_{1},R\cup\sigma_{0}\xrightarrow{}\sigma_{1})$
 Table 2: Symbolic Expressions for the Proposed Actions.
-In this section, we introduce a novel paradigm that integrates the extraction of aspect-opinion pairs into a process reminiscent of constructing a parsing-directed graph.
-This model progressively constructs and annotates aspect-oriented relations using input sequences enriched with contextual features.
-Unlike the classic approach, which employs three sets to record action histories, we utilize five sets, including stack, buffer, aspect, and pair, to show how aspect and opinion are extracted sitimulately.
+Here, $\sigma$ represents the stack, $\beta$ represents the buffer, $A$ denotes the aspect, $O$ denotes the opinion, and $R$ consist of an aspect and an opinion.
+3 The Trans-AOPE/ASTE Model
+In this section, we introduce a novel paradigm that treats the extraction of aspect-opinion pairs as a parsing-directed graph construction process.
+This paradigm comprises two major procedures: aspect-opinion pair extraction and tagging.
+We refer to the pair-extraction model as Trans-AOPE and the model that incorporates tagging as Trans-ASTE.
+Trans-AOPE progressively constructs and annotates aspect-oriented relations using input sequences enriched with contextual features.
+Unlike classic approaches, which employ three sets to record action histories, our paradigm utilizes five sets (including a stack, buffer, aspect set, and pair set) to demonstrate how aspects and opinions are extracted together.
 3.1 Transitional Operations and State Change
@@
 It is important to note that the set of actions shown in the figure is not the only way to extract the "Gourmet food" and "delicious" aspect-opinion pair.
-The alternative solution for parsing the sentence is shown in Appendix 5.
+An alternative approach use the stack’s capacity of holding multiple tokens, moving "is" to the stack ($\beta_{3}\rightarrow\sigma_{3}$) before merging "Gourmet" and "food" ($[\sigma_{1},\sigma_{2},\sigma_{3}]\rightarrow[\sigma_{1\&2},\sigma_{3}]$).
 3.2 Trans-AOPE State Representation
@@
 3.3 Transition Implementation with Neural Model
-This section outlines the transition-based parsing process using neural models.
+This section introduces a transition-based parsing process.
 RoBERTa Liu et al.
-(2019) encodes textual phrases, while UniLSTM Hochreiter and Schmidhuber (1997) and BiLSTM Graves and Schmidhuber (2005) model dynamic transitions.
-The parser state is initialized and evolves through a sequence of actions with distributed embeddings and LSTMs representing the state at each step.
-At the end stage, a Multi-Layer Perceptron (MLP) is employed to classify the sentiment associated with each pair or triplet based on the evolved parser state.
+(2019) encodes the text, while UniLSTM Hochreiter and Schmidhuber (1997) and BiLSTM Graves and Schmidhuber (2005) capture transitions.
+The parser state evolves through a sequence of actions, with LSTMs processing each token once.
+This yields a time complexity of $O(n\cdot d^{2})$, typically simplified to $O(n)$ under fixed $d$.
+Finally, an MLP classifies the sentiment for each pair or triplet based on the final parser state.
 Token representations
@@
 where $N$ is the number of samples, $M$ is the number of classes (either action or sentiment classes), $y_{i}^{j}$ is a binary indicator (0 or 1) indicating whether class $j$ is the correct class for sample $i$, and $p_{i}^{j}$ is the predicted probability for class $j$ for sample $i$.
-The total loss is the sum of the losses for both tasks:
-$\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{sentiment}}.$ (6)
+For AOPE task, total loss is action loss $\mathcal{L}_{\text{action}}$, and for ASTE task, is the sum of the losses for both tasks: $\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{sentiment}}.$
 3.4.2 Contrastive-based Optimization
@@
 where $\omega$ are weighting factors, and $\mathcal{L}_{\text{base}}$ represents any regularization loss added to improve model generalization.
-4 Experiments
-4.1 Dataset
-We used well-established benchmarks in Sentiment Analysis: SemEval 2014 (14lap, 14res) Pontiki et al.
-(2014), SemEval 2015 (15res) Pontiki et al.
-(2015), and SemEval 2016 (16res) Pontiki et al.
+4 Experimental Setups
+4.1 Datasets and preprocessing
+We use well-established benchmarks in Sentiment Analysis: 14lap, 14res from SemEval-2014 Pontiki et al.
+(2014), 15res from SemEval-2015 Pontiki et al.
+(2015), and 16res from SemEval-2016 Pontiki et al.
+(2016) (see Table 4).
+The first is a collection of laptop reviews while the other three are curated from restaurant reviews.
+All are widely used for aspect-based sentiment analysis and extraction tasks Xu et al.
 (2021).
-These datasets, originally curated for the Semantic Evaluation (SemEval) competitions, are widely used for aspect-based sentiment analysis and extraction tasks Xu et al.
-(2021).
-We employed a data processing framework to construct sentiment-enhanced dependency graphs for ASTE datasets.
+Following previous literature, the four datasets are split in a 7:1.5:1.5 proportion for training, development, and test.
+Datasets #S #POS #NEU #NEG #T
+14res 2068 2869 286 754 3909
+14lap 1453 1350 225 774 2349
+15res 1075 1285 61 401 1747
+16res 1393 1674 90 483 2247
+Table 4: Statistics of four datasets. #S denotes the number data pieces, #POS, #NEU, #NEG the number of positive, neutral and negative sentiment labels, and #T the total number of triplets.
+A data processing framework is used to construct sentiment-enhanced dependency graphs for ASTE datasets.
 This framework utilizes SpaCy Honnibal et al.
@@
 Along with tokenized sentences and their corresponding aspect-opinion-sentiment triplets, the processed data supports downstream tasks, including model training and evaluation.
-Figure 3: F1 Score over epochs for the model trained on 14res, 14lap, 15res, and 16res datasets and tested on the 15res dataset.
-The results indicate that training on datasets other than 15res, such as 14res can yield an even better performance, which suggests the presence of richer action patterns in other datasets for the model to learn.
-We evaluated the conformity of our transition operations in extracting aspects and opinions, with action evaluation results listed in Appendix A.3.
-Since our approach predicts actions rather than tokens, we assessed how well the model learns action patterns across individual datasets.
-For instance, as shown in Figure 3, the model did not perform best on the 15res test set when trained on the 15res training set.
-This suggests that other datasets, such as 14res, may contain richer action patterns for the model to learn.
-Consequently, we believe that training on a fused dataset provides a more diverse context, enhancing the model’s ability to learn these actions, which we then evaluated on separate test sets.
-Figure 4: Growth on different weights for AOPE task on 14lap testing set.
-Figure 5: Growth on different weights for ASTE task on 14lap testing set.
+4.2 Training settings
+We experiment with two settings for training data.
+In-domain.
+In this setting, we train with one of the four datasets and test on the test set of the same dataset, e.g., train on 14lap and test on 14lap.
+Combined-train.
+Since our method depends on the model learning the correct action to pair an aspect with an opinion from the training data, we hypothesize that it will have the advantage of being able to make use of training data from diverse domains to learn various actions.
+Thus we experiment with training on two or more training sets combined, to observe whether there is performance gain when more actions are learned.
+Specifically, we train on several training sets together, and evaluate on a single test set.
 4.3 Baselines
-We use two baselines: published scores from prior studies and previous SOTA models run on the fused dataset.
-Regular baselines include classic attention-based methods (e.g., CMLA Wang et al.
-(2017)), pipeline approaches (Peng-Two-stage Peng et al.
-(2020)), sequence-to-sequence frameworks (BARTABSA Yan et al.
-(2021)), sequence-tagging methods (GTS Wu et al.
+We compare our models with two groups of baselines: (1) published results from prior studies that do not have runnable code available and (2) previous State-of-the-Art (SOTA) models re-trained by us using publicly available code base.
+Group (1), as summarized in Table 1, include CMLA Wang et al.
+(2017), Peng-Two-stage Peng et al.
+(2020), BARTABSA Yan et al.
+(2021), GTS Wu et al.
 (2020), JET-BERT Xu et al.
-(2020)), MRC-based models (COM-MRC Zhai et al.
+(2020), COM-MRC Zhai et al.
 (2022), Triple-MRC Zou et al.
-(2024)), graph-based approaches (EMC-GCN Chen et al.
+(2024), EMC-GCN Chen et al.
 (2022), DGCNAP Li et al.
-(2023)), and the SOTA MiniConGTS Sun et al.
+(2023), and the latest MiniConGTS Sun et al.
 (2024).
-In addition to baselines on separate datasets, we selected two state-of-the-art models with publicly available and runnable code: the most recent model, MiniConGTS Sun et al.
-(2024), and the earlier state-of-the-art for pair extraction, BARTABSA Yan et al.
-(2021).
-We used their original configurations and data processing pipelines, modifying only the training data to include all four training sets from 14lap, 14res, 15res, and 16res.
-The models were then separately evaluated on the four corresponding test sets.
-Models 14Res 14Lap 15res 16res
-P R F1 P R F1 P R F1 P R F1
+For re-trained models in group (2), we use two state-of-the-art models with publicly available and runnable code: the most recent model, MiniConGTS Sun et al.
+(2024) [^1] https://github.com/qiaosun22/MiniConGTS, which achieves SOTA performance on both ASTE and AOPE tasks with a complexity of $O(n^{2})$ (fixed constants are omitted for simplicity), and an earlier study, BARTABSA Yan et al.
+(2021) [^2] https://github.com/yhcc/BARTABSA, designed for the AOPE task with a complexity of $O(n)$.
+We use their original configurations and data processing pipelines, modifying only the training data to include all four training sets from 14lap, 14res, 15res, and 16res, when necessary.
+The models are then separately evaluated on the four test sets.
+14res 14lap 15res 16res
+In-domain setting P R F1 P R F1 P R F1 P R F1
 CMLA+Wang et al. (2017) $\diamond$ - - 48.95 - - 44.10 - - 44.60 - - 50.00
@@
 MiniConGTS Sun et al. (2024) $\star$ - - 79.60 - - 73.23 - - 73.87 - - 76.29
-SOTAs on fused dataset
-BARTABSA-fused 75.40 76.76 76.07 72.36 63.40 67.59 88.85 92.77 90.77 87.06 87.32 87.19
-MiniConGTS-fused 81.06 78.82 79.92 75.62 71.61 73.56 93.63 92.72 93.17 78.80 81.82 80.28
-Ours on fused dataset
-Trans-AOPE ($w_{1}$:$w_{2}$ = 1:1) 91.88 80.83 86.00 85.87 74.68 79.89 94.88 88.25 91.45 94.63 89.10 91.80
-Trans-AOPE ($w_{1}$:$w_{2}$ = 1:0) 93.80 80.39 86.58 88.60 74.26 80.80 95.07 88.95 91.91 93.10 89.01 91.01
+Trans-AOPE (Ours) 71.04 67.03 68.98 60.32 44.08 50.94 72.91 66.30 69.45 73.56 64.58 68.78
+Combined-train setting (Training Sets)
+MiniConGTS (14lap & 14res) 75.72 78.20 76.94 71.05 68.64 69.83 63.30 69.90 66.44 68.13 72.54 70.27
+MiniConGTS (14res, 15res, & 16res) 78.68 76.78 77.72 56.86 48.31 52.23 94.54 92.48 93.50 77.65 75.22 76.42
+MiniConGTS (14res, 14lap, 15res, & 16res) 78.22 77.11 77.66 76.32 64.19 69.74 93.35 91.99 92.67 76.27 76.79 76.53
+BARTABSA (14res, 14lap, 15res & 16res) 75.40 76.76 76.07 72.36 63.40 67.59 93.56 94.43 93.56 87.06 87.32 87.19
+Ours
+Trans-AOPE (14lap & 14res) 91.95 83.24 87.38 90.60 79.88 84.91 73.99 73.19 73.59 74.84 69.94 72.31
+Trans-AOPE (14res, 15res & 16res) 74.34 62.29 67.78 91.03 78.11 84.07 95.66 88.04 91.69 90.63 80.65 85.35
+Trans-AOPE (14res, 14lap, 15res & 16res) 92.92 83.61 88.02 92.20 80.47 85.94 96.17 90.94 93.48 93.75 84.82 89.06
 Table 5: Comparison of different models on multiple datasets for AOPE task.
-Recall and precision values are omitted where they are not provided.
-Former best scores are underlined, and current best scores are bold.
+Recall and precision values are omitted where they are not reported.
+The former best scores are underlined, and current best scores are bold.
+Highlights are used for analysis.
 $\diamond$ are retrieved from Yan et al., 2021.
 $\bullet$ is retrieved from Zhao et al., 2020, and $\star$ are retrieved from Sun et al., 2024
-Models 14Res 14Lap 15Res 16Res
-P R F1 P R F1 P R F1 P R F1
+5 Results and Analysis
+In this section, we first investigate the effect of adding contrastive loss on performance to determine the optimal training loss configuration (section 5.1).
+Next, we present the results of experiments under different training data configurations, and compare our model with previous methods (section 5.2).
+We analyze our results in section 5.3.
+5.1 Effect of Contrastive Loss
+To investigate the impact of different weight configurations between the base loss and contrastive loss in Equation 9, we evaluated five ratios in $w_{base}:w_{con}$—–namely $1:0$, $1:1$, $1:10$, $0:1$, and $10:1$—–with a batch size of 4.
+The results on 14lap dataset for the AOPE task are shown in Figure 3.
+Notably, the contrastive-heavy settings where $w_{base}:w_{con}=1:1~{}\text{and}~{}1:10$ achieved stronger early-stage performance, indicating the initial advantages of adding contrastive optimization.
+Figure 3: F1 score as a function of training epochs in the combined-train condition for the AOPE task on the 14lap test set, with various loss weight configurations. w1=base loss; w2=contrastive loss.
+A similar pattern emerged in the ASTE task, where the base and balanced configurations consistently outperformed the other weight settings, and the balanced option exhibited faster improvements than the base setting.
+These trends were also observed across additional datasets.
+Consequently, all subsequent experiments used the balanced configuration $w_{base}:w_{con}=1:1$ for optimization.
+5.2 Main Results
+On AOPE
+First, in the in-domain setting, presented in the upper part of Table 5, when models are trained on one dataset only, we observe that our Trans-model (Trans-AOPE) performs worse than previous models.
+[^3] Note that the numbers for previous models are retrieved from respective papers that describe the system, and we are assuming they were trained with only in-domain training set, since many of them did not report which training set is used.
+Second, for the combined-train condition, Trans-AOPE has an advantage over MiniConGTS in all combinations of training sets.
+What is particularly striking is that, when trained with four training sets combined, Trans-AOPE achieves roughly 20 points or more increase in F1 on all test sets (from 68.98 to 88.02 for 14res, 50.94 to 85.94 for 14lap).
+This is not the case for MiniConGTS, which sees performance drop in two test sets (14res and 14lap), and negligible increase in one (16res), large increase in another (15res: 73.87 to 92.67, but still lower than Trans-AOPE, 93.48).
+This suggests that MiniConGTS is not adaptive and cannot reliably make use of a mixture of training data from two domains and multiple datasets.
+Similar trends are observed for BARTABSA.
+What is more striking is that adding restaurant data to the training set (14res and 14lap) improves performance of Trans-AOPE on the laptop test set, from 50.94 to 84.91, but not for MiniConGTS, which drops from 73.23 to 69.83.
+Moreover, when trained only on the three restaurant datasets (14res, 15res, and 16res), Trans-AOPE achieves performance comparable to what it attains when the 14lap dataset is included in the training set (84.07 in the yellow row, compared to 84.91 in the blue row).
+In contrast, MiniConGTS, trained on those same three restaurant datasets, achieves an F1 score of only 52.23.
+This suggests that Trans-AOPE (unlike previous models) can reliably transfer what it has learned in one dataset to another, even if the data are from a different domain, in the AOPE task.
+14res 14lap 15res 16Res
+In-domain setting P R F1 P R F1 P R F1 P R F1
 Peng-Two-stage Peng et al. (2020) $\star$ 43.24 63.66 51.46 38.87 50.38 42.87 48.07 57.51 52.32 46.96 64.24 54.21
-BARTABSA Yan et al. (2021) $\diamond$ 65.52 64.99 65.25 61.41 56.19 58.69 59.14 59.38 59.26 66.60 68.68 67.62
-GTS Wu et al. (2020) $\star$ 67.76 67.29 67.50 57.82 51.32 54.36 62.59 57.94 60.15 66.08 66.91 67.93
 JET-BERT Xu et al. (2020) $\star$ 70.56 55.94 62.40 55.39 43.57 51.04 64.45 51.96 57.53 70.42 58.37 63.83
 COM-MRC Zhai et al. (2022) $\star$ 75.46 68.91 72.01 58.15 60.17 61.17 68.35 61.24 64.53 71.55 71.59 71.57
-EMC-GCN Chen et al. (2022) $\star$ 71.21 72.39 71.78 61.70 56.26 58.81 61.54 62.47 61.93 65.62 71.30 68.33
-BDTF Chen et al. (2022) $\star$ 75.53 73.24 74.35 68.94 55.97 61.74 68.76 63.71 66.12 71.44 73.13 72.27
 DGCNAP Li et al. (2023) $\star$ 72.90 68.69 70.72 62.02 53.79 57.57 62.23 60.21 61.19 69.75 69.44 69.58
@@
 STAGE-3D Liang et al. (2023) $\star$ 78.58 69.58 73.76 71.98 53.86 61.58 73.63 57.90 64.79 76.67 70.12 73.24
+BARTABSA Yan et al. (2021) $\diamond$ 65.52 64.99 65.25 61.41 56.19 58.69 59.14 59.38 59.26 66.60 68.68 67.62
 MiniConGTS Sun et al. (2024) $\star$ 76.10 75.08 75.59 66.82 60.68 63.61 66.50 63.86 65.15 75.52 74.14 74.83
-SOTAs on fused dataset
-BARTABSA-fused 71.05 72.33 71.68 63.66 56.01 59.59 88.98 92.77 90.77 83.82 84.07 83.95
-MiniConGTS-fused 73.89 72.84 73.36 68.26 57.42 62.37 90.64 89.32 89.98 72.73 73.21 72.97
+Trans-ASTE (Ours) 65.46 59.38 62.27 53.44 41.42 46.67 66.28 61.96 64.04 64.82 59.23 61.90
+Combined-train setting (Training Sets)
+MiniConGTS (14lap & 14res) 72.11 74.48 73.28 62.50 60.38 61.42 56.48 62.38 59.28 64.57 68.75 66.59
+MiniConGTS (14res, 15res, & 16res) 74.64 72.84 73.73 52.12 44.28 47.88 92.31 90.29 91.29 74.42 72.10 73.24
+MiniConGTS (14res, 14lap, 15res & 16res) 73.89 72.84 73.36 68.26 57.42 62.37 90.64 89.32 89.98 72.73 73.21 72.97
+BARTABSA (14res, 14lap, 15res & 16res) 71.05 72.33 71.68 63.66 56.01 59.59 91.30 92.99 92.13 83.82 84.07 83.95
 Ours
-Trans-ASTE ($w_{1}$:$w_{2}$ = 1:1) 89.30 77.10 82.76 83.81 69.64 76.07 93.80 86.71 90.11 92.18 86.81 89.42
-Trans-ASTE ($w_{1}$:$w_{2}$ = 1:0) 91.70 78.24 84.44 84.63 70.59 76.98 94.31 88.11 91.11 92.78 87.58 90.11
+Trans-ASTE (14lap &14res) 88.05 80.51 84.11 88.11 74.56 80.77 69.55 67.03 68.27 64.53 62.80 63.65
+Trans-ASTE (14res, 15res & 16res) 73.53 61.74 67.12 42.25 32.25 36.58 89.89 86.96 88.40 81.19 77.08 79.08
+Trans-ASTE (14res, 14lap, 15res & 16res) 89.56 81.24 85.20 86.87 76.33 81.26 94.34 90.58 92.42 90.03 83.33 86.55
 Table 6: Comparison of different models on multiple datasets for ASTE task.
-Former best scores are underlined, and current best scores are bold.
+Recall and precision values are omitted where they are not reported.
+Highlights are used for analysis.
+The former best scores are underlined, and current best scores are bold.
 $\star$ are retrieved from Sun et al., 2024, and $\diamond$ is retrieved from Yan et al., 2021.
-4.3 Results and Analysis
-4.3.1 Study on Augmented Optimization
-To assess the impact of weight compositions in Equation 10, we tested $(w_{1},w_{2})\in\{(1:0),(1:1),(1:10),(0:1),(10:1)\}$, where $w_{1}$ and $w_{2}$ are the weights for base and contrastive-based optimization, respectively, using a batch size of 4.
-Results for AOPE and ASTE are presented in Figures 5 and 5.
-For AOPE, contrastive-heavy weights ($w_{1}:w_{2}=1:1$ and $w_{1}:w_{2}=1:10$) performed well during early epochs (1–50), demonstrating the early advantages of contrastive optimization.
-After epoch 50, base ($w_{1}:w_{2}=1:0$) and balanced ($w_{1}:w_{2}=1:1$) optimizations provided greater stability and better final performance.
-In ASTE, base and balanced optimizations consistently outperformed others, indicating that ASTE primarily benefits from foundational learning, with limited advantages from contrastive optimization.
-Similar trends were observed across other datasets, as shown in Appendix A.4, highlighting the broader applicability of these findings.
-Although contrastive-enhanced optimization does not deliver the best overall performance, it effectively improves efficiency in pair extraction tasks.
-4.3.2 Comparison with Baselines
-On AOPE
-[…]
 On ASTE
-Results for ASTE are listed in Table 4.
-For the 14Res dataset, Trans-ASTE achieves the highest F1 score (84.44%), representing an APG of 12.76% over BARTABSA-fused.
-Similarly, it outperforms MiniConGTS-fused with an APG of 11.08%.
-In the 14Lap dataset, Trans-ASTE) achieves an APG value of 17.39% in F1 compared to BARTABSA-fused, and 14.61% over MiniConGTS-fused.
-For the 15Res dataset, Trans-ASTE) maintains a high level of performance, with an F1 score (91.11%) slightly leading both baselines.
-The overall F1 improvement is 0.34% compared to BARTABSA-fused and 1.13% compared to MiniConGTS-fused.
-The 16Res dataset also sees consistently high performance from Trans-ASTE), with the highest F1 score (90.11%), representing an APG of 6.16% over BARTABSA-fused.
-Overall, combining datasets for all-for-one training does not always result in performance improvements across all datasets for SOTA models.
-However, when data is fused for our model, it consistently achieves performance gains across datasets, with significant improvements observed in 14Res, 14Lap, and 16Res.
-That said, the slightly lower performance on the 15Res AOPE task highlights potential areas for further optimization in specific scenarios.
-5 Conclusion
-In this paper, we present an efficient transition pipeline for aspect-opinion pair extraction with linear time complexity $O(n)$, enhanced by a contrastive learning-based optimization method.
-This approach eliminates the need to directly identify and extract individual tokens, avoiding token-level bias.
-It can be trained on datasets with the most comprehensive coverage of the designed actions, resulting in significant performance improvements across various datasets.
-Specifically, training our model on a well-covered fused dataset enables it to learn robust action combination patterns, leading to superior performance on all datasets.
-Our model surpasses SOTA models on the same fused dataset, setting new benchmarks for AOPE and ASTE tasks.
-This approach not only enhances model efficiency but also demonstrates its effectiveness and versatility in improving performance.
+From Table 6 we observe similar patterns as AOPE, where our Trans-model demonstrates worse performance in the in-domain setting but superior performance under the combined-train setting.
+In the in-domain setting, Trans-ASTE is on par with previous models in 15res, but behind MiniConGTS on the other three test sets.
+However, under the combined-train condition, we again observe that our Trans-model benefits from the mixture of training data.
+For instance, while testing on 14res, the performance gains a 20 points boost when adding the training data of 14lap to 14res (62.27 to 85.20).
+On the contrary, for MiniConGTS, we see a drop in performance from 75.59 to 73.28.
+For BARTABSA, there is a moderate improvement (65.25 to 71.68).
+When four training sets are combined, Trans-ASTE shows superior performance in all test sets except 15res, usually by a large margin (gray rows in Table 6).
+This is especially true for 14lap, in which Trans-ASTE is the only model benefiting tremendously from adding restaurant data (performance increases from 46.67 to 81.26).
+The only pattern different from Trans-AOPE is that when training on three restaurant datasets and testing on 14lap (yellow row), Trans-ASTE has very low performance of 36.58 F1, much lower than the previous 84.07 F1 in Trans-AOPE.
+We believe that this is due to the inability of the sentiment tagging system to transfer from one domain to another, rather than the issue of the aspect–opinion pair extraction of our transition-based pipeline.
+We thus find that Trans-model demonstrates stronger transfer performance in the AOPE task compared to ASTE, possibly because AOPE only requires predicting actions to extract aspect–opinion pairs.
+In contrast, ASTE requires both identifying these pairs and assigning a sentiment tag to form triplets, and sentiment tags tend to be more domain-specific.
+5.3 When and Why is Trans-model Better?
+When: Trans-model is better with combined training sets and in multi-domain generalization.
+From the results above, it seems clear that the Trans-model in the two tasks are better when trained on multiple datasets combined, rather than one dataset alone.
+Our results also suggest that the added training data do not have to be in the same domain: when trained on 14lap & 14res and tested on either 14lap or 14res, the F1 score is at least about 10 percentage points better than trained on only one of the two datasets.
+It shows that our Trans-model is capable of learning from multiple domains, and seems to be able to transfer its knowledge from the laptop domain to the restaurant domain and vice versa, unlike MiniConGTS, which seems to be more sensitive to the domain of the data.
+However, further research is need to better understand the discrepancy between Trans-AOPE and Trans-ASTE models in cross-domain transfer, and improve the cross-domain transfer ability of the sentiment tagger.
+Why: Trans-model can learn more actions in data from diverse domains.
+We believe that the proposed trans-models excel when trained on multiple datasets and show considerable generalization ability for two main reasons.
+First, by predicting actions instead of tokens, it avoids token-level biases and prevents overfitting on token-level patterns specific to restaurant datasets.
+Additionally, trans-models perform pair extraction after the aspect–opinion relationship is established, which enables the model to capture contextual relationships more effectively.
+These design decisions work together to significantly enhance the overall effectiveness and robustness of the trans-models.
+6 Conclusion and Future Work
+In this paper, we present an efficient transition pipeline for the extraction of aspects-opinion pairs with linear time complexity $O(n)$, enhanced by a contrastive-based optimization method.
+This approach obviates the need to directly identify and extract individual tokens, thereby mitigating token-level bias.
+It can be trained on a combination of diverse datasets that offers the most comprehensive coverage of the actions needed, resulting in significant performance improvements across various datasets.
+Specifically, training our model on a well-covered fused dataset enables it to learn robust action patterns, leading to superior performance on all datasets.
+Our model surpasses retrained baseline models on the same fused dataset, establishing new state-of-the-art results for both AOPE and ASTE tasks.
+As transition-based methods have remained relatively less explored in sentiment-related tasks, we believe our work shows a promising direction to employ such methods in aspect-based sentiment analysis.
+Future work can further examine the potential of transition-based models in other sentiment analysis tasks, as well as the generalization ability of these models in situations of multi-domain data.
+It is also important to better understand the cross-domain and multi-domain generalization ability of transition models, by experimentation on more domains, since only two domains are involved in this work.
 Limitation
-The primary limitation of the transition-based pipeline lies in its reliance on relatively large and diverse datasets to effectively learn action patterns.
-As evidenced by the findings in Figure 3, the model achieved better performance on the 15res test set when trained on datasets like 14res and 16res, which appear to provide richer and more diverse action patterns compared to the 15res training set.
-This observation highlights the inadequacy of relying solely on a single dataset, as it may lack the breadth and complexity needed to capture the full spectrum of action patterns required for robust model generalization.
-Without access to such diverse training data, the model’s ability to adapt to nuanced patterns or unseen contexts in the test data may be significantly compromised, emphasizing the importance of training on a fused or broader dataset to enhance its overall effectiveness.
-Appendix A Appendix
-A.1 Alternative Solution with State Changes
-Since stack can hold more than two tokens, "is" can be moved to stack ($\beta_{3}\rightarrow\sigma_{3}$) before "Gourmet" and "food" merge ($[\sigma_{1},\sigma_{2},\sigma_{3}]\rightarrow[\sigma_{1\&2},\sigma_{3}]$).
-The detailed process is present in Table 5.
-Phrase Action Stack ($\sigma$) Buffer ($\beta$) Aspect Opinion Pair
-[] [$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
-1 $SF$ [$\sigma_{1}$] [$\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
-2 $SF$ [$\sigma_{1}$, $\sigma_{2}$] [$\beta_{3}$, $\beta_{4}$]
-3 $SF$ [$\sigma_{1}$, $\sigma_{2}$, $\beta_{3}$] [$\beta_{4}$]
-4 $M$ [$\sigma_{1\&2}$, $\beta_{3}$] [$\beta_{4}$]
-5 $R_{n}$ [$\sigma_{1\&2}$] [$\beta_{4}$]
-6 $SF$ [$\sigma_{1\&2}$, $\sigma_{4}$] []
-7 $RR$ [$\sigma_{1\&2}$, $\sigma_{4}$] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
-9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
-Table 5: State changes for "Gourmet food is delicious" using symbolic representation.
-Here, $\sigma_{1}$ corresponds to "Gourmet", $\sigma_{2}$ to "food", $\sigma_{3}$ to "is", and $\sigma_{4}$ to "delicious".
-Similarly, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ correspond to tokens in the buffer in sequence.
-A.2 Algorithm for Transition Actions
-In Algorithm 1, while the core of the algorithm focuses on the action operations, the configuration also plays a vital role in its functionality.
-The algorithm begins with a buffer, a list of input tokens representing the text to be parsed.
-These tokens are consumed during the parsing process and form the basis for token pair relationships.
-A critical component of the algorithm is the pair2sent mapping, which associates token pairs with their corresponding sentiment labels, allowing the algorithm to enrich the output with sentiment information, even for token pairs not explicitly provided in the target data (defaulting to a neutral label).
-Then, each parsing action (e.g., Shift, Left Relation, Right Relation) is mapped to a unique numeric ID.
-This mapping enables a compact representation of the sequence of parsing actions, which is useful for downstream tasks like training or evaluation.
-Throughout the parsing process, the algorithm records intermediate states (stack, buffer, actions, and sentiment labels) to maintain a complete history of its operations.
-Simultaneously, valid token pairs are stored in pairs, capturing the structured relationships in the text.
-Algorithm 1 Argument-Opinion Pair Parsing Algorithm
-1:buffer (list of input tokens), pair2sent (mapping of token pairs to sentiment labels), action2id (mapping of parsing actions to numeric IDs)
-2:states (list of recorded parsing states), pairs (list of valid token pairs)
-3:Initialize empty stack, actions, states, and pairs. $\triangleright$ Main processing loop: parse tokens from the buffer
-4:while buffer is not empty do
-5:     if len(stack) < 2 then
-6:         stack.append(buffer.pop(0)) $\triangleright$ Apply $SH$: Shift
-7:         Append $SF$ to actions.
-8:         continue
-9:     end if $\triangleright$ Form a pair from the top two elements in stack
-10:     stackPair $\leftarrow$ (stack[-2], stack[-1])
-11:     currentAction $\leftarrow$ getAction(stackPair, targetPairs)
-12:     Append currentAction to actions. $\triangleright$ Handle relation formation actions ($LR$ and $RR$)
-13:     if currentAction == $LR$ then
-14:         Add stackPair to pairs.
-15:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
-16:         Apply stack[-2] and stack[-1] form a left-relation pair.
-17:     else if currentAction == $RR$ then
-18:         Add stackPair to pairs.
-19:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
-20:         Apply stack[-1] and stack[-2] form a right-relation pair.
-21:     else
-22:         sentimentLabel $\leftarrow 3$triangleright Default sentiment label.
-23:     end if $\triangleright$ Record the current parsing state
-24:     Record (stack, buffer, actions, sentimentLabel) in states. $\triangleright$ Modify the stack based on currentAction
-25:     if currentAction == $L_{n}$ then
-26:         Remove stack[-2] $\triangleright$ Apply $L_{n}$: Left Constituent Removal
-27:     else if currentAction == $R_{n}$ then
-28:         Remove stack[-1] $\triangleright$ Apply $R_{n}$: Right Constituent Removal
-29:     else if currentAction == $M$ then
-30:         Merge stack[-2] and stack[-1] $\triangleright$ Apply $M$: Merge
-31:     end if
-32:end while $\triangleright$ Finalize by reducing the remaining elements in the stack
-33:while len(stack) > 1 do
-34:     Remove the top element from stack $\triangleright$ Apply $ST$: Stop
-35:end while $\triangleright$ Convert actions to numeric IDs for output
-36:Map each action in states to its ID using action2id. return states, pairs.
-A.3 Action Coverage on Different Datasets
-14lap 14res 15res 16res fused
-train (Precision) 99.99 99.99 99.99 99.99 99.99
-train (Recall) 85.51 86.63 92.98 93.04 88.73
-train (F1) 92.19 92.84 96.36 96.40 94.03
-dev (Precision) 99.99 99.99 99.99 99.99 99.99
-dev (Recall) 85.97 86.89 94.92 93.04 89.88
-dev (F1) 92.46 92.99 97.39 96.40 94.67
-test (Precision) 99.99 99.99 99.99 99.99 99.99
-test (Recall) 86.66 87.54 91.47 92.42 89.41
-test (F1) 92.85 93.36 95.54 96.06 94.41
-total (Precision) 99.99 99.99 99.99 99.99 99.99
-total (Recall) 86.02 87.03 92.56 92.79 89.00
-total (F1) 92.48 93.06 96.14 96.26 94.18
-Table 6: Transposed Transitional Action Coverage for 14Lap, 14Res, 15Res, 16Res, and fused Datasets.
-The coverage of actions, as shown in Table 6, reveals some differences in the recall values across data sets.
-The 14Lap and 14Res datasets exhibit lower recall compared to the 15Res and 16Res datasets.
-This trend is observed across training, development, and testing phases, with recall values for 14 datasets remaining around 85-87%, while 15Res and 16Res achieve significantly higher recalls, exceeding 91% in most cases.
-The fused dataset strikes a balance with intermediate recall values, reflecting the integration of characteristics from all datasets.
-This highlights a relative weakness in capturing relevant actions in 14 datasets compared to 15 and 16 datasets.
-A.4 Growth with Different Optimization Weight
-Figures 7 and 7, 9 and 9, and 11 and 11 depict the performance growth with varying weights applied to the 14res, 15res, and 16res datasets, respectively.
-These figures highlight the influence of weight adjustments on model accuracy and efficiency in ASTE tasks.
-By examining the trends across these datasets, we demonstrate the early benefits of augmented optimization in enhancing performance.
-Figure 6: Growth on different weights for AOPE task on 14res testing set.
-Figure 7: Growth on different weights for ASTE task on 14res testing set.
-Figure 8: Growth on different weights for AOPE task on 15res testing set.
-Figure 9: Growth on different weights for ASTE task on 15res testing set.
-Figure 10: Growth on different weights for AOPE task on 16res testing set.
-Figure 11: Growth on different weights for ASTE task on 16res testing set.
-A.5 Growth Comparison for AOPE and ASTE
-The growth comparison for AOPE and ASTE is illustrated in Figures 13, 13, 15, and 15, corresponding to the 14res, 14lap, 15res, and 16res datasets, respectively.
-While ASTE generally follows the trends observed in AOPE, the fluctuations in ASTE are more extreme.
-These figures emphasize the sensitivity of ASTE to variations in model weights or dataset conditions, which results in sharper peaks and deeper troughs compared to the relatively smoother patterns seen in AOPE.
-This highlights the unique challenges posed by the ASTE task across different datasets.
-Figure 12: Growth Comparison for AOPE and ASTE on 14res Testing set.
-Figure 13: Growth Comparison for AOPE and ASTE on 14lap Testing set.
-Figure 14: Growth Comparison for AOPE and ASTE on 15res Testing set.
-Figure 15: Growth Comparison for AOPE and ASTE on 16res Testing set.
+Although the Trans-model demonstrates robust generalization capability, its reliance on larger datasets to effectively learn action patterns remains a notable limitation of the transition-based pipeline.
+This issue is evident in our results: while it is not necessary to use the same dataset for both training and testing, the model performs better when trained on blended datasets rather than on a single, limited one.
+Consequently, if the training data lack sufficient action patterns, the model’s ability to handle nuanced or previously unseen contexts can be significantly compromised.
+These findings underscore the importance of training on a combined or broader dataset to enhance the model’s overall effectiveness.